Parallel light rays with a wavelength of 650 nm fall on a single slit. On a screen 2.90 m away, the distance between the first dark fringes on either side of the central maximum is 4.00 mm .

What is the width of the slit?
Answer #1

The distance between dark fringes gives width of central maxima:

\(\Delta y=\frac{2 \lambda L}{d}\)

\(d=\frac{2 \lambda L}{\Delta y}=\frac{2\left(650 \times 10^{-9}\right)(2.90 \mathrm{~m})}{4.0 \times 10^{-3}}=0.95 \times 10^{-3} \mathrm{~m}=0.95 \mathrm{~mm}\)